# Fine-Tuning Next-Scale Visual Autoregressive Models with Group Relative Policy Optimization

### v1

@@ 2.2 Generative Model Alignment with Reinforcement Learning
Defining what a human considers to be a ‘good’ text sample can be very challenging to do via a loss function. Instead, LLMs are first pre-trained on a vast corpus of internet-scraped text data, minimizing a cross-entropy loss to acquire their base knowledge, and then later aligned through fine-tuning on instructions. Since the introduction of Reinforcement Learning from Human Feedback (RLHF) (Ouyang et al., 2022) to turn base language models into chatbots (e.g. GPT-3 (Brown et al., 2020) into ChatGPT), there has been a plethora of follow up research that has ultimately lead to using RL to teach LLMs how to reason. RLHF leverages a reward model, that is, a neural network that takes in text tokens and returns a scalar value that numerically represents human preference. In this setup, the policy is the language model itself, generating text sequences based on a given prompt. Its action space is the token vocabulary size, and its observation space is all possible input token sequences. Finally, the reward function guides the learning process, combining feedback from the aforementioned reward model with a mechanism to prevent drastic changes to the policy, often based on the Kullback–Leibler (KL) divergence. We will be working with a similar setup in Section 4.
@@ 3 Preliminaries
Traditionally, the next-token prediction paradigm has been adapted for image generation by flattening visual samples, predominantly using a raster-scan approach, where image patches are arranged into a sequence (van den Oord et al., 2017). This disrupts the local spatial correlations in images and other visual data, such as videos, hinders the model’s ability to learn bidirectional correlations within the sequence, and may also negatively affect zero-shot generalization. For instance, it restricts lower-right image patches to depend on upper-left patches but not vice versa. Visual AutoRegressive modeling (VAR) (Tian et al., 2024) suggests using a more natural multi-scale, coarse-to-fine representation for images. In VAR the likelihood of the data sample is modeled as follows: $p([r_{1},r_{2},\dots,r_{K}])=\prod_{k=1}^{K}p(r_{k}\mid[r_{1},r_{2},\dots,r_{k-1}])$, where $r_{k}$ is used to denote an autoregressive unit at each scale $k$ of the visual data representation consisting of $h_{k}\times w_{k}$ tokens, where $h_{k},w_{k}\in\mathbb{N}_{+}$ and $h_{K}>h_{K-1}>...>h_{1}$ and $w_{K}>w_{K-1}>...>w_{1}$. In short, in this formulation we impose a causal autoregressive dependency at the resolution (or scale) level, rather than at the token level. That is, $r_{k}$ is a collection of tokens instead of a single token. This in turn, requires to construct an alternative latent space, where the continuous embedding is represented using quantized multi-scale residual embeddings. Similar to next-token prediction, once the transition probabilities have been learned, it can be used to generate new data samples via the token-to-codebook mappings and passing the final embedding reconstruction to a decoder.
@@ 4.1 Group Relative Policy Optimization (GRPO)
Among the techniques that were recently proposed to fine-tune LMMs with RL, Group Relative Policy Optimization (GRPO) (Shao et al., 2024) has emerged as a particularly effective method, offering improved training efficiency and greater stability by replacing the PPO value function estimator with group sampling. More specifically, given a class label $t$ and a set of $G$ outputs $\{o_{1},o_{2},\ldots,o_{G}\}$ sampled from the policy $\pi_{\theta_{old}}$, the objective of GRPO is to update $\pi_{\theta}$ by maximizing the following function:
$\mathcal{J}(\theta)$ $=\mathbb{E}_{o_{i}\sim\pi_{\theta_{old}}}\Bigg{[}\frac{1}{G}\sum_{i=1}^{G}\min\bigg{(}\frac{\pi_{\theta}(o_{i}\mid t)}{\pi_{\theta_{old}}(o_{i}\mid t)},$ (1)
$\quad\text{clip}\Bigl{(}\frac{\pi_{\theta}(o_{i}\mid t)}{\pi_{\theta_{old}}(o_{i}\mid t)},1-\epsilon,1+\epsilon\Bigr{)}A_{i}\bigg{)}-\beta D_{\text{KL}}\Bigg{]},$
where $\epsilon$ is the clipping threshold hyperparameter (Schulman et al., 2017), $\beta$ controls the strength of the (per token) KL divergence regularization term with respect to the pre-trained policy $\pi_{\theta_{ref}}$ $D_{\text{KL}}(\pi_{\theta}\mid\mid\pi_{\theta_{ref}})=\frac{\pi_{\theta_{ref}}(o_{i}\mid t)}{\pi_{\theta}(o_{i}\mid t)}-\log(\frac{\pi_{\theta_{ref}}(o_{i}\mid t)}{\pi_{\theta}(o_{i}\mid t)})-1$, and $A_{i}$ represents the advantage estimate computed with the rewards obtained within each output group:
@@ 4.2 GRPO for Next-Scale Autoregressive Image Models
It is important to note that during training, we use simple multinomial sampling from the VAR output logits scaled with a temperature value $\tau$ to improve exploration. We deliberately avoid using classifier-free guidance (CFG) (Ho & Salimans, 2022), top-p or top-k sampling, beam search, or other techniques that could improve sample quality in order to maintain a tractable sampling distribution during optimization. During inference, however, we do use CFG along with top-p or top-k sampling. More concretly, we employ the same hyperparameters as those used in the original VAR model (Tian et al., 2024).

### v2

@@ 2.2 Generative Model Alignment with Reinforcement Learning
Defining what a human considers to be a ‘good’ text sample can be very challenging to do via a loss function. Instead, LLMs are first pre-trained on a vast corpus of internet-scraped text data, minimizing a cross-entropy loss to acquire their base knowledge, and then later aligned through fine-tuning on instructions. Since the introduction of Reinforcement Learning from Human Feedback (RLHF) (Ouyang et al., 2022) to turn base language models into chatbots (e.g. GPT-3 (Brown et al., 2020) into ChatGPT), there has been a plethora of follow-up research that has ultimately lead to using RL to teach LLMs how to reason. RLHF leverages a reward model, that is, a neural network that takes in text tokens and returns a scalar value that numerically represents human preference. In this setup, the policy is the language model itself, generating text sequences based on a given prompt. Its action space is the token vocabulary size, and its observation space is all possible input token sequences. Finally, the reward function guides the learning process, combining feedback from the aforementioned reward model with a mechanism to prevent drastic changes to the policy, often based on the Kullback–Leibler (KL) divergence. We will be working with a similar setup in Section 4.
@@ 3 Preliminaries
Traditionally, the next-token prediction paradigm has been adapted for image generation by flattening visual samples, predominantly using a raster-scan approach, where image patches are arranged into a sequence (van den Oord et al., 2017). This disrupts the local spatial correlations in images and other visual data, such as videos, hinders the model’s ability to learn bidirectional correlations within the sequence, and may also negatively affect zero-shot generalization. For instance, it restricts lower-right image patches to depend on upper-left patches but not vice versa. Visual AutoRegressive modeling (VAR) (Tian et al., 2024) suggests using a more natural multi-scale, coarse-to-fine representation for images. In VAR the likelihood of the data sample is modeled as follows: $p([r_{1},r_{2},\dots,r_{K}])=\prod_{k=1}^{K}p(r_{k}\mid[r_{1},r_{2},\dots,r_{k-1}])$, where $r_{k}$ is used to denote an autoregressive unit at each scale $k$ of the visual data representation consisting of $h_{k}\times w_{k}$ tokens, where $h_{k},w_{k}\in\mathbb{N}_{+}$ and $h_{K}>h_{K-1}>...>h_{1}$ and $w_{K}>w_{K-1}>...>w_{1}$. In short, in this formulation we impose a causal autoregressive dependency at the resolution (or scale) level, rather than at the token level. That is, $r_{k}$ is a collection of tokens instead of a single token. This, in turn, requires to construct an alternative latent space, where the continuous embedding is represented using quantized multi-scale residual embeddings. Similar to next-token prediction, once the transition probabilities have been learned, it can be used to generate new data samples via the token-to-codebook mappings and passing the final embedding reconstruction to a decoder.
@@ 4.1 Group Relative Policy Optimization (GRPO)
Among the techniques that were recently proposed to fine-tune LMMs with RL, Group Relative Policy Optimization (GRPO) (Shao et al., 2024) has emerged as a particularly effective method, offering improved training efficiency and greater stability by replacing the PPO value function estimator with group sampling. More specifically, given a class label $c$ and a set of $G$ outputs $\{o_{1},o_{2},\ldots,o_{G}\}$ sampled from the policy $\pi_{\theta_{old}}$, the objective of GRPO is to update $\pi_{\theta}$ by maximizing the following function:
$\mathcal{J}(\theta)$ $=\mathbb{E}_{o_{i}\sim\pi_{\theta_{old}}}\Bigg{[}\frac{1}{G}\sum_{i=1}^{G}\min\bigg{(}\frac{\pi_{\theta}(o_{i}\mid c)}{\pi_{\theta_{old}}(o_{i}\mid c)},$ (1)
$\quad\text{clip}\Bigl{(}\frac{\pi_{\theta}(o_{i}\mid c)}{\pi_{\theta_{old}}(o_{i}\mid c)},1-\epsilon,1+\epsilon\Bigr{)}A_{i}\bigg{)}-\beta D_{\text{KL}}\Bigg{]},$
where $\epsilon$ is the clipping threshold hyperparameter (Schulman et al., 2017), $\beta$ controls the strength of the (per token) KL divergence regularization term with respect to the pre-trained policy $\pi_{\theta_{ref}}$ $D_{\text{KL}}(\pi_{\theta}\mid\mid\pi_{\theta_{ref}})=\frac{\pi_{\theta_{ref}}(o_{i}\mid c)}{\pi_{\theta}(o_{i}\mid c)}-\log(\frac{\pi_{\theta_{ref}}(o_{i}\mid c)}{\pi_{\theta}(o_{i}\mid c)})-1$, and $A_{i}$ represents the advantage estimate computed with the rewards obtained within each output group:
@@ 4.2 GRPO for Next-Scale Autoregressive Image Models
It is important to note that during training, we use simple multinomial sampling from the VAR output logits scaled with a temperature value $\tau$ of 0.7 to balance exploration and exploitation. We deliberately avoid using classifier-free guidance (CFG) (Ho & Salimans, 2022), top-p or top-k sampling, beam search, or other techniques that could improve sample quality in order to maintain a tractable sampling distribution during optimization. During inference, however, we do use CFG along with top-p or top-k sampling. More concretly, we employ the same hyperparameters as those used in the original VAR model (Tian et al., 2024).